OpenStudy (anonymous):

express in terms of i (a+bi form) -√(-144)=

OpenStudy (anonymous):

-(12i)=-12i = 0-12i

OpenStudy (anonymous):

sqrt-1=i, sqrt144=12

jimthompson5910 (jim_thompson5910):

-√(-144)= -√(-1*12^2)=-√(-1)√(12^2)=-i*12 = -12i So the answer is -12i